# Improving Sequential Market Coordination via Value-oriented Renewable Energy Forecasting

### v2

@@ IV-A The Impact of the Forecast on the DA and RT Solutions
This section analyzes the impact of the forecasts on the primal and dual solutions of DA and RT market clearings. Specifically, we derive analytical functions that quantitatively depict the impact of the forecasts on the optimal primal and dual solutions. As a parameter to the DA market clearing (10), forecasts $\hat{\bm{y}}_{d}$ influence DA primal and dual solutions directly. As for the RT market clearing, the impact of the forecast is indirect, as it does not directly appear as the parameter in (11) or (12). Concretely, for the RT market clearing at time-slot $\tau=1$, the forecast $\hat{\bm{y}}_{d}$ influences $\bm{x}_{d,1}^{*}$, and then $\bm{x}_{d,1}^{*}$ influences the RT solutions, with the first impact being determined by the DA clearing (10), and the second impact by the RT clearing (11). For the RT clearing at time-slot $\tau=2,...,T$, the impact of the forecast is more complex. The forecast $\hat{\bm{y}}_{d}$ affects the DA solutions $\bm{x}_{d,\tau}^{*},\bm{p}_{d,\tau-1}^{*}$ through the DA clearing (10). The influence of the forecast $\hat{\bm{y}}_{d}$ on the RT solutions $\bm{p}_{d,\tau-1}^{+-*}$ at the previous time-slot $\tau-1$ occurs as explained earlier. Then, the influence of the parameters $\bm{x}_{d,\tau}^{*},\bm{p}^{*}_{d,\tau-1},\bm{p}^{+-*}_{d,\tau-1}$ on the RT clearing solutions at time $\tau$ is determined by (12). The above process is summarized in Fig. 3.
@@
[22] Consider the linear program (16) and the parameter $\bm{\omega}\in\Omega$. The primal and dual decision policies are a piecewise linear function and a stepwise function respectively, if there exists a polyhedral partition $R_{1},...,R_{N}$ of $\Omega$, and $\forall\bm{\omega}\in R_{i}$, the primal decision policy is linear, and the dual decision policy is a constant function.
@@
With Proposition 1, we present the local primal decision policy for the DA clearing (10), and RT clearing (11) and (12). Let $\mathcal{J}^{a}_{\text{DA},d},\mathcal{J}^{a}_{\text{RT},d,\tau},\forall\tau=1,...,T$ denote the row index set of active constraints (10b), (11b), and (12b). The parameters associated with the active constraints are denoted as $\bm{G}_{\text{DA}}[\mathcal{J}_{\text{DA},d}^{a}],\bm{\psi}_{\text{DA},d}[\mathcal{J}_{\text{DA},d}^{a}],\bm{F}_{\text{DA}}^{y}[\mathcal{J}_{\text{DA},d}^{a}]$, $\bm{G}_{\text{RT}}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{\psi}_{\text{RT},d,\tau}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{F}_{\text{RT}}[\mathcal{J}_{\text{RT},d,\tau}^{a}]$,$\bm{G}_{\text{RT}}^{\prime}[\mathcal{J}_{\text{RT},d,\tau}^{a}]$, and $\bm{\psi}_{\text{RT},d,\tau}^{\prime}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{F}_{\text{RT}}^{\prime x}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{F}_{\text{RT}}^{\prime p}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{F}_{\text{RT}}^{\prime+-}[\mathcal{J}_{\text{RT},d,\tau}^{a}]$. We have the following proposition for the local primal decision policies of (10), (11), and (12).

### v3

@@ IV-A The Impact of the Forecast on the DA and RT Solutions
This section analyzes the impact of the forecasts on the primal and dual solutions of DA and RT market clearings. Specifically, we derive analytical functions that quantitatively depict the impact of the forecasts on the optimal primal and dual solutions, whose illustration is shown in Fig. 2. As a parameter to the DA market clearing (10), forecasts $\hat{\bm{y}}_{d}$ influence DA primal and dual solutions directly. As for the RT market clearing, the impact of the forecast is indirect, as it does not directly appear as the parameter in (11) or (12). Concretely, for the RT market clearing at time-slot $\tau=1$, the forecast $\hat{\bm{y}}_{d}$ influences $\bm{x}_{d,1}^{*}$, and then $\bm{x}_{d,1}^{*}$ influences the RT solutions, with the first impact being determined by the DA clearing (10), and the second impact by the RT clearing (11). For the RT clearing at time-slot $\tau=2,...,T$, the impact of the forecast is more complex. The forecast $\hat{\bm{y}}_{d}$ affects the DA solutions $\bm{x}_{d,\tau}^{*},\bm{p}_{d,\tau-1}^{*}$ through the DA clearing (10). The influence of the forecast $\hat{\bm{y}}_{d}$ on the RT solutions $\bm{p}_{d,\tau-1}^{+-*}$ at the previous time-slot $\tau-1$ occurs as explained earlier. Then, the influence of the parameters $\bm{x}_{d,\tau}^{*},\bm{p}^{*}_{d,\tau-1},\bm{p}^{+-*}_{d,\tau-1}$ on the RT clearing solutions at time $\tau$ is determined by (12).
@@
([22]) Consider the linear program (16) and the parameter $\bm{\omega}\in\Omega$. The primal and dual decision policies are a piecewise linear function and a stepwise function respectively, if there exists a polyhedral partition $R_{1},...,R_{N}$ of $\Omega$, and $\forall\bm{\omega}\in R_{i}$, the primal decision policy is linear, and the dual decision policy is a constant function.
@@
With Proposition 1, we present the local primal decision policy for the DA clearing (10), and RT clearing (11) and (12). Let $\mathcal{J}^{a}_{\text{DA},d},\mathcal{J}^{a}_{\text{RT},d,\tau},\forall\tau=1,...,T$ denote the row index set of active constraints (10b), (11b), and (12b). The parameters associated with the active constraints are denoted as $\bm{G}_{\text{DA}}[\mathcal{J}_{\text{DA},d}^{a}],\bm{\psi}_{\text{DA},d}[\mathcal{J}_{\text{DA},d}^{a}],\bm{F}_{\text{DA}}^{y}[\mathcal{J}_{\text{DA},d}^{a}]$, $\bm{G}_{\text{RT}}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{\psi}_{\text{RT},d,\tau}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{F}_{\text{RT}}[\mathcal{J}_{\text{RT},d,\tau}^{a}]$$\bm{G}_{\text{RT}}^{\prime}[\mathcal{J}_{\text{RT},d,\tau}^{a}]$, and $\bm{\psi}_{\text{RT},d,\tau}^{\prime}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{F}_{\text{RT}}^{\prime x}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{F}_{\text{RT}}^{\prime p}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{F}_{\text{RT}}^{\prime+-}[\mathcal{J}_{\text{RT},d,\tau}^{a}]$. We have the following proposition for the local primal decision policies of (10), (11), and (12).